OpenStudy (anonymous):

The sum of three consecutive odd integers is –15. What are the numbers?

OpenStudy (anonymous):

You call your number x

OpenStudy (ghazi):

x+x+2+x+4= -15 3x+6=-15 3x=-21 x=-7

OpenStudy (ghazi):

so the numbers are -7, -7+2, -7+4 -7,-5,-3

OpenStudy (anonymous):

could it also be 7 9 11 all negative

OpenStudy (anonymous):

-7, -9, -11 will not give you the sum of -15.. -7 -9 - 11 = -27 (\(\ne 15\))